OpenStudy (anonymous):

Choose the Standard Form of the equation y = −one over fourx − 6. one over fourx + y = −6 4x + y = −6 x + 4y = −24 −4x + y = −24

OpenStudy (shayaan_mustafa):

one over fourx + y =-6 ---or--- y=-one over fourx - 6 can be written as x/4 + y = -6 (x + 4y)/4 = -6 x + 4y = -24 (Second one is correct.)